OpenStudy (richyw):

Find the global maximum and the global minimum of the function \(f(x,y)=e^{-x^2-y^2}(x^2+2y^2)\) on the plate region R consisting of all points for which \(x^2+y^2\leq 4\) by finding the critical points of f(x,y) inside R and on the boundary of R. You can use the fact that e>2

OpenStudy (richyw):

So I think my first order of buisness is to find the \(\nabla f(x,y)\) which I get \[\left\langle2xe^{-x^2-y^2}(1-x^2-2y^2),2ye^{-x^2-y^2}(2-x^2-2y^2)\right\rangle\]

OpenStudy (richyw):

now would I set this equal to \[\left\langle 0,0\right\rangle\]

OpenStudy (richyw):

giving me \[2xe^{-x^2-y^2}(1-x^2-2y^2)=0\]\[2ye^{-x^2-y^2}(2-x^2-2y^2)=0\]

OpenStudy (richyw):

Is this correct so far? If so I can see that (0,0) would be a critical point. so would I just need to solve \[1-x^2-2y^2=0\] and \[2-x^2-2y^2=0\]

OpenStudy (richyw):

No I am really lost here. No Idea what to do.

OpenStudy (turingtest):

well one solution you seem to have overooked is \[x=0,y=0\]

OpenStudy (turingtest):

overlooked*

OpenStudy (richyw):

I mentioned that actually. I can't figure out the other ones!

OpenStudy (turingtest):

you mean you want to look at the boundary?

OpenStudy (turingtest):

I keep getting zero for everything... not sure if that's right

OpenStudy (richyw):

I'm not sure if I'm doing anything correct haha.

OpenStudy (turingtest):

I am sort of familiar with these problems, but am a bit rusty here is an example of a... somewhat similar problem http://tutorial.math.lamar.edu/Classes/CalcIII/AbsoluteExtrema.aspx (example 2) but this is harder and more confusing than the example given

OpenStudy (anonymous):

didn't think this is calculus

OpenStudy (turingtest):

and what did you think it was panlac?

OpenStudy (anonymous):

precal

OpenStudy (anonymous):

Is there a guide for the equation tool within the website? I see sometimes the text are huge and there are many things that aren't included in the tools.